OpenStudy (anonymous):

A data set consists of the following data points: (2, 4), (4, 7), (5, 12) The line of best fit has the equation = -1.5 + 2.5x. What does this equation predict for a value of x = 7? A. 20.5 B. 17.5 C. 16 D. 19

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

hello

Directrix (directrix):

y = -1.5 + 2.5x Evaluate the function for x = 7 y = -1.5 + 2.5 * 7 y = ? @2kende

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

wut do u mean evaluate the function?

Directrix (directrix):

Multiply 2.5 times 7 and post what you get, okay?

OpenStudy (anonymous):

17.5

Directrix (directrix):

Add - 1.5 to 17.5 and post what you get, okay?

OpenStudy (anonymous):

16

Directrix (directrix):

What does this equation predict for a value of x = 7? y = -1.5 + 2.5x y = -1.5 + 2.5 * 7 y = 16 Correct.

OpenStudy (anonymous):

<3

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

im still kinda stuck. can u help wit another?